# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

### 09 Mar 2026 v1

@@ 3.3 Weighted Bregman divergence and information-geometric identities
$\rho^{\mathrm{w}}_{\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\int_{\cal X}\varphi(x)\,p_{\theta_{1}}(x)^{\alpha}p_{\theta_{2}}(x)^{1-\alpha}\,\mathrm{d}\mu(x)$ (3.17)
$=\exp\{\hat{F}(\theta_{\alpha})-\alpha F(\theta_{1})-(1-\alpha)F(\theta_{2})\}.$
@@
$D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2})-\hat{F}(\theta_{\alpha})$ (3.18)
@@
In particular, when $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and the normalised geometric mixture $p_{\theta_{1}}^{\alpha}p_{\theta_{2}}^{1-\alpha}$ belongs to the same exponential family, namely $\propto p_{\theta_{\alpha}}$ with $\theta_{\alpha}=\alpha\theta_{1}+(1-\alpha)\theta_{2}$.
@@
Assume that $\hat{F}$ is strictly convex on the segment $\theta_{1},\theta_{2}]$ and that the maximiser $\alpha^{*}\in(0,1)$ exists. Then $\alpha^{*}$ is unique and satisfies
$(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})=F(\theta_{1})-F(\theta_{2}),\qquad\theta_{\alpha^{*}}=\alpha^{*}\theta_{1}+(1-\alpha^{*})\theta_{2},$ (3.19)
@@
Differentiate (3.18) with respect to $\alpha$ and use $\frac{\mathrm{d}}{\mathrm{d}\alpha}\theta_{\alpha}=\theta_{1}-\theta_{2}$. Strict convexity of $\hat{F}$ on $\theta_{1},\theta_{2}]$ implies strict concavity of $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$, hence uniqueness. ∎
@@
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})=D^{\mathrm{w}}_{B,\alpha^{*}}(p_{\theta_{1}},p_{\theta_{2}})=\alpha^{*}F(\theta_{1})+(1-\alpha^{*})F(\theta_{2})-\hat{F}(\theta_{\alpha^{*}}).$ (3.20)
@@
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}}),$ (3.21)
@@
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})$ $=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}})$ (3.22)
$=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}}).$ (3.23)
In the special case $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and $E_{\varphi}(\theta)\equiv 1$, so that (3.20) reduces to the classical Jensen divergence induced by $F$ and (3.22) becomes $D_{C}(p_{\theta_{1}},p_{\theta_{2}})=B_{F}(\theta_{1},\theta_{\alpha^{*}})=B_{F}(\theta_{2},\theta_{\alpha^{*}})$.
@@
Equation (3.20) is (3.18) at $\alpha=\alpha^{*}$. For (3.21), expand $B^{\mathrm{w}}_{\varphi,F}(\theta_{i},\theta_{\alpha^{*}})=E_{\varphi}(\theta_{\alpha^{*}})\big[F(\theta_{i})-F(\theta_{\alpha^{*}})-(\theta_{i}-\theta_{\alpha^{*}})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})\big]$ and use (3.19) to see that the difference vanishes. Finally, substituting $\theta_{1}-\theta_{\alpha^{*}}=(1-\alpha^{*})(\theta_{1}-\theta_{2})$ into $B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})/E_{\varphi}(\theta_{\alpha^{*}})$ and using (3.19) gives
@@
Since $\hat{F}(\theta_{\alpha^{*}})=F(\theta_{\alpha^{*}})+\ln E_{\varphi}(\theta_{\alpha^{*}})$, this yields (3.22). ∎
Derivative and weighted KL.
Recall $F_{pq}(\alpha)=\ln Z_{pq}(\alpha)=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$. Assuming differentiation under the integral sign is justified, we have
$F_{pq}^{\prime}(\alpha)=\mathbb{E}_{(pq)_{\alpha}}\!\left[\ln\frac{p(X)}{q(X)}\right],$ (3.24)
where $(pq)_{\alpha}$ is the Chernoff-tilted density from (3.5). In particular,
$F_{pq}^{\prime}(1)=\frac{1}{E_{\varphi}(p)}\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x)=\frac{1}{E_{\varphi}(p)}D^{\mathrm{w}}_{\rm KL}(p\|q).$ (3.25)
@@
$D^{\mathrm{w}}_{C}(p,q)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{1}\big)-\ln E_{\varphi}(p)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{0}\big)-\ln E_{\varphi}(q),$ (3.26)
@@
Here $D_{\rm KL}$ denotes the standard (unweighted) Kullback–Leibler divergence.
@@
At $\alpha=\alpha^{*}$ we have $F^{\prime}_{pq}(\alpha^{*})=0$. Since $F_{pq}(1)=\ln E_{\varphi}(p)$ and $F_{pq}(0)=\ln E_{\varphi}(q)$, the claim follows from $D_{C}^{\mathrm{w}}(p,q)=-F_{pq}(\alpha^{*})$. ∎
@@
Assume that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior, [It is always inferior because of convexity?] so that $F^{\prime}_{pq}(\alpha^{*})=0$. Then
$D_{C}^{\mathrm{w}}(p,q)=B_{F_{pq}}(1,\alpha^{*})-\ln E_{\varphi}(p)=B_{F_{pq}}(0,\alpha^{*})-\ln E_{\varphi}(q).$ (3.27)
@@
and similarly for 0. Substituting this into Lemma 3.12 yields (3.27). ∎
@@
Assume $d=1$, $\theta_{1}\neq\theta_{2}$, and that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior. Assume moreover that $\hat{F}^{\prime}$ is strictly increasing on $\Theta$ and set ${\hat{G}}=(\hat{F}^{\prime})^{-1}$. Then (3.19) yields
$\alpha^{*}=\frac{1}{\theta_{1}-\theta_{2}}\left({\hat{G}}\left(\frac{F(\theta_{1})-F(\theta_{2})}{\theta_{1}-\theta_{2}}\right)-\theta_{2}\right).$ (3.28)
When $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and (3.28) reduces to the classical formula.
To illustrate the general identities above, we provide explicit expressions for $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ in several parametric settings (Gaussian, Poisson, and exponential), see Section 4.
@@ 3.4 Tilted weighted likelihood and concentration bounds
Although the optimal decision rule in the context-sensitive setting is still based on the usual likelihood ratio $q/p$ (cf. Section 3), it is convenient to consider the tilted likelihood ratio $q^{*}/p^{*}$ as well. The reason is that the context weight $\varphi$ cancels pointwise in $q^{*}/p^{*}$ and affects inference only through the normalisation constants $E_{\varphi}(p)$ and $E_{\varphi}(q)$. We record below (i) a large-deviation representation for $L^{*}/n$ via the cumulant generating functions $\psi_{\mathbb{P}},\psi_{\mathbb{Q}}$ and their Legendre transforms, and (ii) a finite-$n$ concentration bound for $L^{*}$ based on standard martingale inequalities (Azuma–Hoeffding).
@@
$\begin{array}[]{l}L^{*}({X}_{1}^{n})=L^{*}(X_{1},\ldots,X_{n})=\sum\limits_{i=1}^{n}\ln\frac{q^{*}(X_{i})}{p^{*}(X_{i})}\\ \hskip 38.41139pt=\sum\limits_{i=1}^{n}\ln\frac{q(X_{i})}{p(X_{i})}-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p).\end{array}$ (3.29)
@@
Thus $L^{*}$ is the usual log-likelihood ratio, shifted by a constant determined by the context weight $\varphi$.
@@
$=\ln\int_{\cal X}q(x)^{\alpha}p(x)^{1-\alpha}\,\mathrm{d}\mu(x)-\alpha\ln E_{\varphi}(q)+\alpha\ln E_{\varphi}(p),$ (3.30)
@@
$\psi_{\mathbb{Q}}(\alpha):$ $=\ln\mathbb{E}_{\mathbb{Q}}\left[e^{\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]=\ln\mathbb{E}_{\mathbb{P}}\left[e^{\ln\frac{q(X)}{p(X)}+\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]$ (3.31)
@@
$I_{\mathbb{Q}}(r)=I_{\mathbb{P}}(r)-r+\ln E_{\varphi}(p)-\ln E_{\varphi}(q).$ (3.32)
@@
$=:{\hat{D}}^{\rm w}_{C}(q,p).$ (3.33)
Interpretation. The value $I_{\mathbb{P}}(0)$ is the Chernoff–Cramér exponent controlling the tail event $\{L^{*}/n\geq 0\}$ under $\mathbb{P}$, i.e. (under standard regularity assumptions) $\mathbb{P}(L^{*}(X_{1}^{n})\geq 0)\asymp e^{-nI_{\mathbb{P}}(0)}$. In the unweighted case $\varphi\equiv 1$ we have $E_{\varphi}(p)=E_{\varphi}(q)=1$ and $I_{\mathbb{P}}(0)$ reduces to the classical Chernoff information between $p$ and $q$. We also stress that ${\hat{D}}_{C}^{\rm w}(q,p)=I_{\mathbb{P}}(0)$ is a tilted-likelihood exponent and is distinct from the weighted Chernoff information $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$ from Definition 2.4, which governs the optimal sum of context-sensitive losses.
Non-asymptotic concentration via a Doob martingale.
@@
$=\sum\limits_{j=1}^{k}\ln\frac{q(X_{j})}{p(X_{j})}+(n-k)D_{\rm KL}(\mathbb{Q}\|\mathbb{P})-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p),$ (3.34)
@@
$U_{k}-U_{k-1}=\ln\frac{q(X_{k})}{p(X_{k})}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P}),\qquad k=1,\ldots,n.$ (3.35)
@@
On the next step we consider a particular case when $d<\infty$ and $\sigma^{2}<\infty$, where $d=\sup\limits_{x\in{\cal X}}\left|\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right|$ and
@@
$=\int_{\cal X}q(x)\left(\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right)^{2}\,\mathrm{d}\mu(x).$ (3.36)
@@
$\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\},$ (3.37)
@@
We use the following modified version of Azuma–Hoeffding inequality, see [2].
@@
$\begin{array}[]{l}|U_{k}-U_{k-1}|\leq d,\\ {\rm Var}[U_{k}-U_{k-1}|{\cal F}_{k-1}]\leq\sigma^{2}.\end{array}$ (3.38)
@@
${\mathbb{P}}\left(|U_{n}-U_{0}|\geq\beta n\right)\leq 2\exp\left\{-nD\left(\frac{\delta+\gamma}{1+\gamma}\bigg\|\frac{\gamma}{1+\gamma}\right)\right\},$ (3.39)
@@
${\mathbb{P}}_{\mathbb{Q}}\big(L^{*}(X_{1},\ldots,X_{n})\geq\beta n\big)\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\}.$ (3.40)
Lemma 3.15 is quoted from [2]. Theorem 3.16 is its direct application to the tilted log-likelihood ratio (3.29); the only dependence on the context weight $\varphi$ is through $E_{\varphi}(p)$ and $E_{\varphi}(q)$.

### 09 May 2026 v3

@@ 3.3.2 Weighted Chernoff/Bhattacharyya Quantities Inside an Exponential Family
$\rho^{\mathrm{w}}_{\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\ln\int_{\cal X}\varphi(x)\,p_{\theta_{1}}(x)^{\alpha}p_{\theta_{2}}(x)^{1-\alpha}\,\mathrm{d}\mu(x)$ (30)
$=\hat{F}(\theta_{\alpha})-\alpha F(\theta_{1})-(1-\alpha)F(\theta_{2}).$
@@
$D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2})-\hat{F}(\theta_{\alpha})$ (31)
@@
In particular, when $\varphi\equiv 1$, we have $\hat{F}\equiv F$ and the normalised geometric mixture $p_{\theta_{1}}^{\alpha}p_{\theta_{2}}^{1-\alpha}$ belongs to the same exponential family, namely, $\propto p_{\theta_{\alpha}}$ with $\theta_{\alpha}=\alpha\theta_{1}+(1-\alpha)\theta_{2}$.
@@
Assume that $\hat{F}$ is strictly convex on the segment $\theta_{1},\theta_{2}]$ and that the maximiser $\alpha^{*}\in(0,1)$ exists. Then, $\alpha^{*}$ is unique and satisfies
$(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})=F(\theta_{1})-F(\theta_{2}),\qquad\theta_{\alpha^{*}}=\alpha^{*}\theta_{1}+(1-\alpha^{*})\theta_{2},$ (32)
@@
Differentiate (31) with respect to $\alpha$ and use $\frac{\mathrm{d}}{\mathrm{d}\alpha}\theta_{\alpha}=\theta_{1}-\theta_{2}$. Strict convexity of $\hat{F}$ on $\theta_{1},\theta_{2}]$ implies strict concavity of $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$, hence uniqueness. ∎
@@
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})=D^{\mathrm{w}}_{B,\alpha^{*}}(p_{\theta_{1}},p_{\theta_{2}})=\alpha^{*}F(\theta_{1})+(1-\alpha^{*})F(\theta_{2})-\hat{F}(\theta_{\alpha^{*}}).$ (33)
@@
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}}),$ (34)
@@
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})$ $=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}})$ (35)
$=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}}).$ (36)
In the special case $\varphi\equiv 1$, we have $\hat{F}\equiv F$ and $E_{\varphi}(\theta)\equiv 1$, so that (33) reduces to the classical Jensen divergence induced by $F$ and (35) becomes $D_{C}(p_{\theta_{1}},p_{\theta_{2}})=B_{F}(\theta_{1},\theta_{\alpha^{*}})=B_{F}(\theta_{2},\theta_{\alpha^{*}})$.
@@
Equation (33) is (31) at $\alpha=\alpha^{*}$. For (34), expand $B^{\mathrm{w}}_{\varphi,F}(\theta_{i},\theta_{\alpha^{*}})=E_{\varphi}(\theta_{\alpha^{*}})\big[F(\theta_{i})-F(\theta_{\alpha^{*}})-(\theta_{i}-\theta_{\alpha^{*}})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})\big]$ and use (32) to see that the difference vanishes. Finally, substituting $\theta_{1}-\theta_{\alpha^{*}}=(1-\alpha^{*})(\theta_{1}-\theta_{2})$ into $B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})/E_{\varphi}(\theta_{\alpha^{*}})$ and using (32) gives
@@
Since $\hat{F}(\theta_{\alpha^{*}})=F(\theta_{\alpha^{*}})+\ln E_{\varphi}(\theta_{\alpha^{*}})$, this yields (35). ∎
3.3.3 Derivative and Weighted KL
Recall $F_{pq}(\alpha)=\ln Z_{pq}(\alpha)=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$. In view of (14 ), the differentiation under the integral sign is justified, and we have
$F_{pq}^{\prime}(\alpha)=\mathbb{E}_{(pq)_{\alpha}}\!\left[\ln\frac{p(X)}{q(X)}\right],$ (37)
where $(pq)_{\alpha}$ is the Chernoff-tilted density from (18). In particular,
$F_{pq}^{\prime}(1)=\frac{1}{E_{\varphi}(p)}\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x)=\frac{1}{E_{\varphi}(p)}D^{\mathrm{w}}_{\rm KL}(p\|q).$ (38)
@@ 3.3.3 Derivative and Weighted KL
$D^{\mathrm{w}}_{C}(p,q)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{1}\big)-\ln E_{\varphi}(p)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{0}\big)-\ln E_{\varphi}(q),$ (39)
@@
Here, $D_{\rm KL}$ denotes the standard (unweighted) Kullback–Leibler divergence.
@@
At $\alpha=\alpha^{*}$, we have $F^{\prime}_{pq}(\alpha^{*})=0$. Since $F_{pq}(1)=\ln E_{\varphi}(p)$ and $F_{pq}(0)=\ln E_{\varphi}(q)$, the claim follows from $D_{C}^{\mathrm{w}}(p,q)=-F_{pq}(\alpha^{*})$. ∎
@@
Assume that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior, so that $F^{\prime}_{pq}(\alpha^{*})=0$. Then
$D_{C}^{\mathrm{w}}(p,q)=B_{F_{pq}}(1,\alpha^{*})-\ln E_{\varphi}(p)=B_{F_{pq}}(0,\alpha^{*})-\ln E_{\varphi}(q).$ (40)
@@
and similarly for 0. Substituting this into Lemma 3.11 yields (40). ∎
@@
Assume $d=1$, $\theta_{1}\neq\theta_{2}$, and that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior. Assume moreover that $\hat{F}^{\prime}$ is strictly increasing on $\Theta$ and set ${\hat{G}}=(\hat{F}^{\prime})^{-1}$. Then (32) yields
$\alpha^{*}=\frac{1}{\theta_{1}-\theta_{2}}\left({\hat{G}}\left(\frac{F(\theta_{1})-F(\theta_{2})}{\theta_{1}-\theta_{2}}\right)-\theta_{2}\right).$ (41)
When $\varphi\equiv 1$, we have $\hat{F}\equiv F$ and (41) reduces to the classical formula.
To illustrate the general identities above, we provide explicit expressions for $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ in several parametric settings (Gaussian, Poisson, and exponential); see Section 4.
Remark 3.14.
The Bregman representation (40) identifies $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$ with a Bregman divergence on the tilted log-normaliser $\hat{F}$; geometrically, the optimiser $\alpha^{\ast}$ marks the intersection of the exponential geodesic of the tilted family $\{\varphi\,p^{\alpha}q^{1-\alpha}\}_{\alpha\in[0,1]}$ with the weighted Kullback–Leibler bisector, generalising the unweighted characterisation of [11, 12]. In the exponential-family setting, the computation of $D_{C}^{\mathrm{w}}$ therefore reduces to $\hat{F}$ and its gradient: once $\hat{F}$ is available in closed form, $\alpha^{\ast}$ is determined by (41) in the one-parameter case and by a monotone equation in the natural-parameter space in general, without evaluating $\rho_{\alpha}^{\mathrm{w}}$ for each $\alpha$. The examples of Section 4 exemplify this reduction.
@@ 3.4 Tilted Weighted Likelihood and Concentration Bounds
Although the optimal rule in the context-sensitive problem is still the usual likelihood-ratio test $q/p$ (cf. Section 3), it is convenient to work with the tilted ratio $q^{*}/p^{*}$. The factor $\varphi$ cancels pointwise in $q^{*}/p^{*}$ and enters only through the normalisation constants $E_{\varphi}(p)$ and $E_{\varphi}(q)$. We record two consequences: a large-deviation representation for $L^{*}/n$ via the cumulant generating functions $\psi_{\mathbb{P}},\psi_{\mathbb{Q}}$ and their Legendre transforms and a finite-$n$ concentration bound based on a martingale argument.
@@
$\begin{array}[]{l}L^{*}({X}_{1}^{n})=L^{*}(X_{1},\ldots,X_{n})=\sum\limits_{i=1}^{n}\ln\frac{q^{*}(X_{i})}{p^{*}(X_{i})}\\ \hskip 38.41139pt=\sum\limits_{i=1}^{n}\ln\frac{q(X_{i})}{p(X_{i})}-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p).\end{array}$ (42)
@@
Thus, $L^{*}$ is the usual log-likelihood ratio, shifted by a constant determined by the context weight $\varphi$.
@@
$=\ln\int_{\cal X}q(x)^{\alpha}p(x)^{1-\alpha}\,\mathrm{d}\mu(x)-\alpha\ln E_{\varphi}(q)+\alpha\ln E_{\varphi}(p),$ (43)
@@
$\psi_{\mathbb{Q}}(\alpha):$ $=\ln\mathbb{E}_{\mathbb{Q}}\left[e^{\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]=\ln\mathbb{E}_{\mathbb{P}}\left[e^{\ln\frac{q(X)}{p(X)}+\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]$ (44)
@@
$I_{\mathbb{Q}}(r)=I_{\mathbb{P}}(r)-r+\ln E_{\varphi}(p)-\ln E_{\varphi}(q).$ (45)
@@
$=:{\hat{D}}^{\rm w}_{C}(q,p).$ (46)
Interpretation. The value $I_{\mathbb{P}}(0)$ is the Chernoff–Cramér exponent controlling the tail event $\{L^{*}/n\geq 0\}$ under $\mathbb{P}$, i.e., (under standard regularity assumptions), $\mathbb{P}(L^{*}(X_{1}^{n})\geq 0)\asymp e^{-nI_{\mathbb{P}}(0)}$. In the unweighted case $\varphi\equiv 1$, we have $E_{\varphi}(p)=E_{\varphi}(q)=1$ and $I_{\mathbb{P}}(0)$ reduces to the classical Chernoff information between $p$ and $q$. We also stress that ${\hat{D}}_{C}^{\rm w}(q,p)=I_{\mathbb{P}}(0)$ is a tilted-likelihood exponent and is distinct from the weighted Chernoff information $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$ from Definition 2.4, which governs the optimal sum of context-sensitive losses.
Non-Asymptotic Concentration Via a Doob Martingale
@@
$=\sum\limits_{j=1}^{k}\ln\frac{q(X_{j})}{p(X_{j})}+(n-k)D_{\rm KL}(\mathbb{Q}\|\mathbb{P})-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p),$ (47)
@@
$U_{k}-U_{k-1}=\ln\frac{q(X_{k})}{p(X_{k})}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P}),\qquad k=1,\ldots,n.$ (48)
@@
Assume now that $d<\infty$ and $\sigma^{2}<\infty$, where $d=\sup\limits_{x\in{\cal X}}\left|\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right|$ and
@@
$=\int_{\cal X}q(x)\left(\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right)^{2}\,\mathrm{d}\mu(x).$ (49)
@@
$\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\},$ (50)
@@
We use the following modified version of Azuma–Hoeffding inequality; see [14].
@@
$\begin{array}[]{l}|U_{k}-U_{k-1}|\leq d,\\ {\rm Var}[U_{k}-U_{k-1}|{\cal F}_{k-1}]\leq\sigma^{2}.\end{array}$ (51)
@@
${\mathbb{P}}\left(|U_{n}-U_{0}|\geq\beta n\right)\leq 2\exp\left\{-nD\left(\frac{\delta+\gamma}{1+\gamma}\bigg\|\frac{\gamma}{1+\gamma}\right)\right\},$ (52)
@@
${\mathbb{P}}_{\mathbb{Q}}\big(L^{*}(X_{1},\ldots,X_{n})\geq\beta n\big)\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\}.$ (53)
Lemma 3.15 is quoted from [14]. Theorem 3.16 is its direct application to the tilted log-likelihood ratio (42); the only dependence on the context weight $\varphi$ is through $E_{\varphi}(p)$ and $E_{\varphi}(q)$.